OpenStudy (anonymous):

Solve the following equation. Then place the correct number in the box provided. 8 + x = 13 + 2x

OpenStudy (anonymous):

Okay, so try to subtract x from each side to where you get x by itself on the right side of the equation.

OpenStudy (anonymous):

After you've done that, tell us what you got.

OpenStudy (anonymous):

So it's 7?

OpenStudy (anonymous):

You need to find what x is equal to, so in this case you would do this: 8 + x = 13 + 2x -x -x 8 = 13 + x

OpenStudy (anonymous):

You now need to get x by itself, so what would you do to get that?

OpenStudy (anonymous):

I'm lost

OpenStudy (anonymous):

All right, it's fine. Everyone always gets lost, but that's how they learn :) The equation is 8 + x = 13 + 2x and you need to get the "x" by itself so it would equal a number.

OpenStudy (anonymous):

To get "x" by itself, you need to subtract the "x" from the side with the least amount of x's. That would get you "x" alone on the right side. Also, if you didn't know as a tip, x = 1x; they're the same thing.

OpenStudy (anonymous):

So it's not 7? I won't graduate if I don't finish this :(

OpenStudy (anonymous):

Don't stress, it's normal to not understand something quickly. :) How did you get 7?